# Layered Multiple Scattering Approach to Hard X-ray Photoelectron Diffraction: Theory and Application

### v1

@@ 2.1 Brief review of PED simulation progress
Basically, there exist three classifications for adequate PED simulation for solids as presented in Fig. 1: cluster-based models, layer-by-layer approaches, and the so-called ”lattice-plane” methods.
Figure 1: Schematic representation of (a) cluster-based, (b) layer-by-layer and (c) lattice-plane models. Atoms are represented by blue. Localized sources are indicated by red and surrounded by spherical waves. Black/blue arrows illustrate single-site/multiple scattering paths.
The first group (Fig. 1a) is a right call for short-range probes (e.g. PED or Auger electron diffraction) and broken crystal symmetry (for instance by disorders). Because of the electron’s short inelastic scattering length, scattering can be limited to a cluster of individual atoms surrounded by the photo-emitter. A model can be formulated by either single-scattering cluster (SSC) [40, 41, 42, 43, 44, 45] or multiple-scattering cluster (MSC) [13, 46] methods. For comparison in this subset, the most precise interpretation is MSC with spherical waves [14]. Specifically, the multiple-scattering effects have to be taken into account while investigating single-crystal because the XPD patterns originate from photoemission of ten (and more) surface layers. Inspired by the work of Kaduwela and co-workers [14] where multiple scattering (MS) is derived from the Rehr-Albers (R-A) separable propagator approximation [47] and applied for initial and final states, more CD measurements are nicely reproduced [48, 49]. Furthermore, the experiment-theory match is improved by utilizing an accurate representation of the Green’s function propagator for evaluating MS expansion [24, 50]. Thereafter, more indispensable modifications and implementations come into existence as seen in following MSC packages: MSCD [16], PAD2 [51], MSPHD [52], EDAC [24] and MsSpec [18]. Recently, Rehr et al. [53] has discussed the development of real-space Green’s function approach for PED based on the R-A formalism and boosted the importance of sharing theoretical software among groups. Workflow tools such as AiiDA [54] and Corvus [55] were proposed to enable advanced and efficient computation of PED without requiring significant changes to the original code.
@@
On the other hand, the second class (Fig. 1b) considers a crystal as a collection of identical layers and makes use of the translational symmetry along surfaces, which is beneficial for describing interlayer multiple scattering in a finite-depth crystal or a surface formed by inequivalent layers on a substrate of repeat periodicity. This category is developed on the basis of low energy electron diffraction (LEED) [72, 73, 74, 75]. Recently, the one-step model (OSM) of photoemission [76] is proposed to conduct a nicely theoretical prediction of the core-level problem of W(110) [33] where the final-state wave function is represented by a so-called time-reversed LEED-state. A 50% difference in the spin-orbit doublets of W(110) was reported in the work of employing circularly polarized radiation. Herein, the sub-levels of 3d and 4p created by SOC splitting (as well as possibly via multiplet splitting and circular polarization) have distinct diffractograms [35]. This generalized angle-resolved core-level photoemission takes into account the fully relativistic effects and experimental geometry. OSM is applicable not only in reciprocal space but also in real space. Arising from the above-mentioned cluster models, Krüger et al. [77] suggested a real-space multiple scattering method integrating OSM and claimed good-agreement VB-PED for Cu(111), which is comparable to the result of k-space calculations. Lately, a reasonable comparison has been made between very low energy electron diffraction (VLEED), CL-PED, and VB-PED for monolayer graphene by the variational embedding method including OSM [78] from the team of Krasovskii. Even though the specifics may be somewhat diverse, diffraction patterns overall display comparable broad elements. Based on LEED theory, it turns out that more hybrid schemes reveal from simple to sophisticated modifications to elucidate complex surface structures. Deriving from tensor LEED (a perturbation technique to high-speed computation of LEED intensity) introduced by Rous et al. [79], PED intensity is able to be computed with a little adjustment to how photoelectron emitter atomic displacements are handled [80]. Combined with Helmholtz’s reciprocity theorem, which was initially stated for optics, the path-reversed LEED was effectively employed for PED [81]. Next, it was extended to crystal surfaces with many atoms per unit cell, leading to the fact that the outcomes are almost identical to those from traditional forward-path computations [82].
@@
When the photoelectron kinetic energies fall into the range of several keV, the hard X-ray photoemission diffraction signal is dominated by scattering in the full, three-dimensional, periodicity of the crystal structure, which results in pronounced bulk features and vanishing sensitivity to surface effects. Tailored to this regime, a third simulation group (Fig. 1c) is based on the dynamical theory of high energy electron diffraction, adapted to the case of photoelectron sources in a single crystal. The approach presented in [21] uses the Bloch wave approach to electron diffraction [83], with a plane wave expansion of the diffracted electron waves in a bulk crystal [84]. The basic building blocks of the theoretical Bloch wave simulation model are the Fourier coefficients of the real and absorptive parts of the 3D crystal potential with corresponding reciprocal lattice vectors, and it can be conceptually advantageous that these theoretical entities are in very direct correspondence to experimentally observed Kikuchi band features [30]. In the context of Kikuchi-band theory, high kinetic energy PED patterns from single crystals can thus be meaningfully discussed in terms of photoelectron reflections at lattice planes, with a Kikuchi band having a width of twice the corresponding Bragg angle [85, 86, 87]. Examples of core-level XPD in the hard X-ray range interpreted by Bloch wave simulations can be found in [22, 32].
@@ 2.2 Theoretical model
The initial state is described by the core-level Green matrix:
$G_{1}(E)=\sum_{\mathcal{K}_{m}}\left\lvert\mathcal{K}_{m}\right\rangle\frac{\Sigma_{c}}{\left(E-\varepsilon_{\mathcal{K}_{m}}\right)^{2}+\Sigma_{c}^{2}}\left\langle\mathcal{K}_{m}\right\rvert,$ (3)
@@
with $E_{fi}=-2\text{i}c/[(E_{f}+c^{2})^{2}-(E_{i}+c^{2})^{2}]$. $E_{i}$ and $E_{f}$ are the energy of initial and final states, respectively. The derivation of the expression involves applying commutator and anticommutator rules in a similar manner to the nonrelativistic case described in [94]. Herein, $\mathbf{A_{0}}$ is the spatially constant amplitude of the electromagnetic vector potential. $V(\mathbf{r})$ and $\mathbf{B}(r)$ refer to the potential and magnetic field, respectively.
@@
On the way to the sample surface, the outgoing electron wave interacts with the surrounding of emitters. To gain a qualitative understanding of diffraction patterns, it is crucial to distinguish the two diffraction regimes based on the relatively medium range of the kinetic energy of photoelectrons. In typical energies less than about $\sim$300 eV, backscattering is strong [97, 1] and it can be utilized in various ways to gather structural data about atoms that are positioned ”behind” the emitter when observed by the detector [14]. When the kinetic energy is $\sim$500 eV or higher, the forward scattering (zeroth-order scattering) mode is dominant [1]. A while back, Thompson and Fadley [98] conducted theoretical simulations to contrast the XPD at 1 and 10 keV for the emission of C 1s from a CO molecule oriented vertically. This means that sharper forward scattering diffraction characteristics are typically anticipated at higher energies. Nevertheless, the higher-order diffraction characteristics are weaker. Due to the significant kinetic energy, forward scattering is predominant in the cases under our current discussion. This treatment is sketched via the inset in Fig. 1b where single-site forward scattering is employed within each layer. Between layers, multiple scattering is taken into account. To obtain better visualization, we make use of surface atoms from the side view. Instead of the many scattering paths (short blue arrows) that need to be treated in low-energy calculations, the forward-scattering paths (long blue arrows) strongly prevail. In addition, two examples for different reciprocal lattice vectors $\vec{G}_{hkl}$ illustrate schematically the analog to the model of Bragg reflection at lattice planes. Under this regime, the diffraction pattern is mainly characterized by features resembling ”Kikuchi patterns”, such as lines and bands typically seen in high-energy electron diffraction. In order to favorably replicate the results from our experiments, we utilized at least 137 $\vec{G}_{hkl}$ for this expansion in the calculations presented here. This approach yielded results that numerically converged with respect to the Kikuchi diffraction pattern. The angular-momentum cutoff, a second convergence parameter of our calculations, necessitates terminating orbital angular momenta at a specific value due to constraints on computing time and memory. We conducted convergence experiments up to $l_{max}=11$ and observed that for $l_{max}>3$, the positions of the photoemission peaks and, importantly for the current study, the CDAD patterns remained qualitatively unchanged. We verified this convergence test using MsSpec [18], a real-space cluster approach. Undoubtedly, we are aware that we would need to approach $l_{max}>30$ (as calculated from the height of centrifugal barrier [99]) in order to make a quantitative comparison between experimental and theoretical intensities for the high kinetic energies of electrons examined here.
@@ 2.3 Experiment
Progress on the electron detector side towards 2D and 3D recording has been accompanied by advances in soft and hard X-ray synchrotron beamlines and X-ray free electron lasers. At the beamline P22 of PETRA III a total energy resolution (combined resolution of ToF analyzer and photon bandwidth) of 62 meV at 6 keV (resolving power $10^{5}$) was achieved [107]. At the free-electron laser FLASH (DESY, Hamburg) a similar ToF momentum microscope is operated [h] and the first time-resolved XPS and XPD experiments with fs resolution at soft X-ray energies have been performed [108, 109, 110]. These technical advances on both the source and detector side have led to an enormous increase in the information content since the early days when CDAD was discovered with rotating electron spectrometers [111, 112], confirming theoretical models of Ritchie [113], Cherepkov and Mc Koy et al. [114, 115].

### v3

@@ 4.1 Brief review of PED simulation progress
Basically, there exist three classifications for adequate PED simulation for solids as presented in Fig. LABEL:fig:Fig2-1.jpg: cluster-based models, layer-by-layer approaches, and the so-called ”lattice-plane” methods.
The first group (Fig. LABEL:fig:Fig2-1.jpga) is a right call for short-range probes (e.g. PED or Auger electron diffraction) and broken crystal symmetry (for instance by disorders). Because of the electron’s short inelastic scattering length, scattering can be limited to a cluster of individual atoms surrounded by the photo-emitter. A model can be formulated by either single-scattering cluster (SSC) [60, 61, 62, 63, 64, 65, 66] or multiple-scattering cluster (MSC) [13, 67, 68] methods. For comparison in this subset, the most precise interpretation is MSC with spherical waves [14]. Specifically, the multiple-scattering effects have to be taken into account while investigating single-crystal because the XPD patterns originate from photoemission of ten (and more) surface layers. Inspired by the work of Kaduwela and co-workers [14] where multiple scattering (MS) is derived from the Rehr-Albers (R-A) separable propagator approximation [69] and applied for initial and final states, more CD measurements are nicely reproduced [70, 71]. Furthermore, the experiment-theory match is improved by utilizing an accurate representation of the Green’s function propagator for evaluating MS expansion [24, 72]. Thereafter, more indispensable modifications and implementations come into existence as seen in following MSC packages: MSCD [16], PAD2 [73], MSPHD [74], EDAC [24] and MsSpec [18]. Recently, Rehr et al. [75] has discussed the development of real-space Green’s function approach for PED based on the R-A formalism and boosted the importance of sharing theoretical software among groups. Workflow tools such as AiiDA [76] and Corvus [77] were proposed to enable advanced and efficient computation of PED without requiring significant changes to the original code.
@@
On the other hand, the second class (Fig. LABEL:fig:Fig2-1.jpgb) considers a crystal as a collection of identical layers and makes use of the translational symmetry along surfaces, which is beneficial for describing interlayer multiple scattering in a finite-depth crystal or a surface formed by inequivalent layers on a substrate of repeat periodicity. This category is developed on the basis of low energy electron diffraction (LEED) [93, 94, 95, 96]. Recently, the one-step model (OSM) of photoemission [57] is proposed to conduct a nicely theoretical prediction of the core-level problem of W(110) [33] where the final-state wave function is represented by a so-called time-reversed LEED-state. A 50% difference in the spin-orbit doublets of W(110) was reported in the work of employing circularly polarized radiation. Herein, the sub-levels of 3d and 4p created by SOC splitting (as well as possibly via multiplet splitting and circular polarization) have distinct diffractograms [35]. This generalized angle-resolved core-level photoemission takes into account the fully relativistic effects and experimental geometry. OSM is applicable not only in reciprocal space but also in real space. Arising from the above-mentioned cluster models, Krüger et al. [97] suggested a real-space multiple scattering method integrating OSM and claimed good-agreement VB-PED for Cu(111), which is comparable to the result of k-space calculations. Lately, a reasonable comparison has been made between very low energy electron diffraction (VLEED), CL-PED, and VB-PED for monolayer graphene by the variational embedding method including OSM [98] from the team of Krasovskii. Even though the specifics may be somewhat diverse, diffraction patterns overall display comparable broad elements. Based on LEED theory, it turns out that more hybrid schemes reveal from simple to sophisticated modifications to elucidate complex surface structures. Deriving from tensor LEED (a perturbation technique to high-speed computation of LEED intensity) introduced by Rous et al. [99], PED intensity is able to be computed with a little adjustment to how photoelectron emitter atomic displacements are handled [100]. Combined with Helmholtz’s reciprocity theorem, which was initially stated for optics, the path-reversed LEED was effectively employed for PED [101]. Next, it was extended to crystal surfaces with many atoms per unit cell, leading to the fact that the outcomes are almost identical to those from traditional forward-path computations [102].
@@
When the photoelectron kinetic energies fall into the range of several keV, the hard X-ray photoemission diffraction signal is dominated by scattering in the full, three-dimensional, periodicity of the crystal structure, which results in pronounced bulk features and vanishing sensitivity to surface effects. Tailored to this regime, a third simulation group (Fig. LABEL:fig:Fig2-1.jpgc) is based on the dynamical theory of high energy electron diffraction, adapted to the case of photoelectron sources in a single crystal. The approach presented in [21] uses the Bloch wave approach to electron diffraction [103], with a plane wave expansion of the diffracted electron waves in a bulk crystal [104]. The basic building blocks of the theoretical Bloch wave simulation model are the Fourier coefficients of the real and absorptive parts of the 3D crystal potential with corresponding reciprocal lattice vectors, and it can be conceptually advantageous that these theoretical entities are in very direct correspondence to experimentally observed Kikuchi band features [30]. In the context of Kikuchi-band theory, high kinetic energy PED patterns from single crystals can thus be meaningfully discussed in terms of photoelectron reflections at lattice planes, with a Kikuchi band having a width of twice the corresponding Bragg angle [105, 106, 107]. Examples of core-level XPD in the hard X-ray range interpreted by Bloch wave simulations can be found in Refs. [22, 32].
@@ 4.2 Theoretical model
The initial state is described by the imaginary part of the core-level Green matrix:
$-\frac{1}{\pi}ImG_{1}(E)=\sum_{\mathcal{K}_{m}}\left\lvert\mathcal{K}_{m}\right\rangle\frac{\Sigma_{c}}{\left(E-\varepsilon_{\mathcal{K}_{m}}\right)^{2}+\Sigma_{c}^{2}}\left\langle\mathcal{K}_{m}\right\rvert,$ (3)
@@
with $E_{fi}=-2\text{i}c/[(E_{f}+c^{2})^{2}-(E_{i}+c^{2})^{2}]$. $E_{i}$ and $E_{f}$ are the energy of initial and final states, respectively. The derivation of the expression involves applying commutator and anticommutator rules in a similar manner to the nonrelativistic case described in [114]. Herein, $\mathbf{A_{0}}$ is the spatially constant amplitude of the electromagnetic vector potential. The scalar potential $V(\mathbf{r})$ and the magnetic exchange field $\mathbf{B}(r)$ are obtained by self-consistent electronic structure calculations performed within the ab-initio framework of spin-density functional theory by use of the Vosko et al. [115] parameterization of the exchange and correlation potential. To account for, among others, impurity scattering a small constant imaginary value of $V_{0i}=0.01$ eV was used for the initial core state. For the final state, a constant imaginary value of $V_{0i}=1$ eV has been chosen again in a phenomenological way.
@@
On the way to the sample surface, the outgoing electron wave interacts with the surrounding of emitters. To gain a qualitative understanding of diffraction patterns, it is crucial to distinguish the two diffraction regimes based on the relatively medium range of the kinetic energy of photoelectrons. In typical energies less than about $\sim$300 eV, backscattering is strong [118, 1] and it can be utilized in various ways to gather structural data about atoms that are positioned ”behind” the emitter when observed by the detector [14]. When the kinetic energy is $\sim$500 eV or higher, the forward scattering (zeroth-order scattering) mode is dominant [1]. A while back, Thompson and Fadley [119] conducted theoretical simulations to contrast the XPD at 1 and 10 keV for the emission of C 1s from a CO molecule oriented vertically. This means that sharper forward scattering diffraction characteristics are typically anticipated at higher energies. Nevertheless, the higher-order diffraction characteristics are weaker. Due to the significant kinetic energy, forward scattering is predominant in the cases under our current discussion. This treatment is sketched via the inset in Fig. LABEL:fig:Fig2-1.jpgb where single-site forward scattering is employed within each layer. Between layers, multiple scattering is taken into account. To obtain better visualization, we make use of surface atoms from the side view. Instead of the many scattering paths (short blue arrows) that need to be treated in low-energy calculations, the forward-scattering paths (long blue arrows) strongly prevail. In addition, two examples for different reciprocal lattice vectors $\vec{G}_{hkl}$ illustrate schematically the analog to the model of Bragg reflection at lattice planes. Under this regime, the diffraction pattern is mainly characterized by features resembling ”Kikuchi patterns”, such as lines and bands typically seen in high-energy electron diffraction. In order to favorably replicate the results from our experiments, we utilized at least 137 $\vec{G}_{hkl}$ for this expansion in the calculations presented here. This approach yielded results that numerically converged with respect to the Kikuchi diffraction pattern. The angular-momentum cutoff, a second convergence parameter of our calculations, necessitates terminating orbital angular momenta at a specific value due to constraints on computing time and memory. We conducted convergence experiments up to $l_{max}=11$ and observed that for $l_{max}>3$, the positions of the photoemission peaks and, importantly for the current study, the CDAD patterns remained qualitatively unchanged. We verified this convergence test using MsSpec [18], a real-space cluster approach. Undoubtedly, we are aware that we would need to approach $l_{max}>30$ (as calculated from the height of centrifugal barrier [120]) in order to make a quantitative comparison between experimental and theoretical intensities for the high kinetic energies of electrons examined here.
@@ 4.3 Experiment
Progress on the electron detector side towards 2D and 3D recording has been accompanied by advances in soft and hard X-ray synchrotron beamlines and X-ray free electron lasers. At the beamline P22 of PETRA III a total energy resolution (combined resolution of ToF analyzer and photon bandwidth) of 62 meV at 6 keV (resolving power $10^{5}$) was achieved [128]. At the free-electron laser FLASH (DESY, Hamburg) a similar ToF momentum microscope is operated [129] and the first time-resolved XPS and XPD experiments with fs resolution at soft X-ray energies have been performed [130, 131, 132]. These technical advances on both the source and detector side have led to an enormous increase in the information content since the early days when CDAD was discovered with rotating electron spectrometers [133, 134], confirming theoretical models of Ritchie [135], Cherepkov and Mc Koy et al. [136, 137].